OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = (x – 5); (–4, 3) a) y = –3x + 15 b) y = –3x – 9 c} y = –4x + 3 d} y = –1/3x – 11

OpenStudy (anonymous):

from the equation y + 2 = (x – 5); it is clear that its slope is 1 hence the slope of the line perpendicular to it is 1/-1=-1 and since it passes through (-4,3) the eq of the reqd line is of the form (y-3)=-1(x+4)

OpenStudy (anonymous):

can u re -write ur question